OpenStudy (anonymous):

Combine like terms. 9m – 12m

OpenStudy (solomonzelman):

\[9m – 12m~~~~~->~~~~~(9-12)m~~~~~~~->~~~~~~(-3)m~~~~->~~~~-3m\]

OpenStudy (solomonzelman):

Hope this helps.

OpenStudy (anonymous):

can u help me with like 3 more

OpenStudy (solomonzelman):

Yes, sure.

OpenStudy (anonymous):

Evaluate the expression, using the given value of the variable. 4 – 2x + 5x when x = 4

OpenStudy (solomonzelman):

\[4-2x+5x~~~~~~~~~~~~~~~~~~~~~~~~~~x=2\]\[4-2(2)+5(2)\]\[4-4+10\]so what is the final answer?

OpenStudy (the_fizicx99):

Plug in for x 4 - 2(4) + 5(2) 4 - 8 + 10

OpenStudy (anonymous):

10??

OpenStudy (solomonzelman):

Sorry.

OpenStudy (solomonzelman):

I got it wrong, you plug in 4, not 2.

OpenStudy (anonymous):

so 14

OpenStudy (the_fizicx99):

Me?? You!!!

OpenStudy (anonymous):

@tHe_FiZiCx99 so 14???

OpenStudy (solomonzelman):

yes

OpenStudy (the_fizicx99):

Yes

OpenStudy (anonymous):

Solve for x. 3 • 4 – 3 • 7 = 3(x – 7)

OpenStudy (the_fizicx99):

3 * 4 - 3*7 = 3x -21 12 - 21 = 3x -21 -9 = 3x - 21 +21 +21 12 = 3x /3 /3 Solve?